clear fractions or decimals, solve, and check.
0.7(3x + 6) = 1.1 - (x - 3)

Answer by checkley77(12844)

0.7(3x + 6) = 1.1 - (x - 3)
2.1x+4.2=1.1-x+3
2.1x+x=4.1-4.2
3.1x=-.1
x=-.1/3.1 or -.032258 ans.
Proof:
.7(3*-.032258+6)=1.1-(-.032258-3)
.7(-.096774+6)=1.1-(-3.032258)
.7*5.903226=1.1+3.032258
4.132258=4.132258